(10 pts) Solve the system S 21 +22-3.23=-2 | 3x1+4x2–3x3= 2
Answer #1

Que :- solve the system J 36, +36 - 3x3 = -2 1 34 +48 - 3X3 = 2 We home hace 4 + x - 3x3 = -2 34 + 4X9-3X=2 In Matrix form LThen x = 8-65 Then by cis, we have x+ 8-65 - 35= -2 3) 24 +8-95 - - 2) 4 = -2-8 +99 4= -10 +95 Therefore 1 + - LoL I